Part A and B on both please. Will upvote.
A conducting sphere of radius rı 0.22 m Problem 7: a23 m has a total charge of e-2-pC. A second uncharged conducting (10%) sphere of radius to their size. links to the first by a thin conducting wire. The spheres are separated by an enormous distance compared 50% Part (a) what is the total charge on sphere two 02 after they are connected, in coulom Grade S Deductio Potential sin() | cos() | tan() |I (IM 7| 8|9 cotan ain acos0 atan acota sinh0 cosh tanh c Submiss Attempts (4% per detailed cotanhO Degrees O Radians Submit Hnt Feedbac k: % deduction per feedback. Hints:-0% deduction per hint. Hints remaining- - -& 50% Part (b) what is the surface charge density of the second sphere, 02, after they are connected in coulombs per square meter?
(10% Problem 6 An electrostatic paint spray erhasa 0.19 m ground that repels paint droplets onto a grounded object. an ter metal sphere at a potential of25.0 kV with respect to the 50% Part (a) what charge is on the sphere, in coulombs? Grade Summary Deductions Potential 100% 0% Submissions Aftempts remaining (4% per attenpo cotan asin acos atan acotan) sinbo cosh tanh0 cotanh0 Degrees O Radians ailed vievw I give up Subnit Hint Feedback: 0% deduction per feedback Hints: 0% deduction Per hint Hints remaining 2. in coulombs, must a 0.105 mg drop o f paint have in order to arrive at the object with a speed of 10.0 ms, starting 50% Part (b) what charge from rest?
Answer #1

charge on q1 = 2.7 micro coulombs

charge on q2 = 0

when they are connected net charge = ( q1 + q 2 ) / 2

                                                                  = 2.7 / 2 = 1.35

charge on q2 = 1.35 micro coulombs

b) surface charge density = q2 / area = q2 / 4 pie r ^2   = 1.35.5 x 10^-6 / 4 * 3.14 * 0.22^2 = 2.2 x 10^-6 c/m^2

2) radius = r = 0.19/2 = 0.095 m

potential = V = 25 x 10^3 V

V = k q / r

charge = q = v r / k = 25 x 10^3 * 0.095 / 9 x 10^9 = 0.264 X 10^-6 = 2.64 x 10^-7 C

b) mass = m = 0.105 x 10^-6 k g

initial speed = V0 = 0 m/sec

final speed = V = 10 m/sec

from work energy theorem W = 1/ 2 m ( V^2 - V0^2 )

                                                  V q = 1/2 m V^2

                                                25000 * q = 1/2 * 0.105 x 10^-6 * 10^2

                                                   charge = q = 2.1 x 10^-10 C